# Enhancing Diffusion-Based Sampling with Molecular Collective Variables

### v1

@@ 4.2 Sampling Reactive Energy Landscapes
While peptide results highlight the efficiency of WT-ASBS in terms of the number of energy evaluations, the computational budgets in diffusion-based sampling in these cases are dominated by SDE integration with a control network, since classical force fields are extremely fast to evaluate. However, classical force fields are insufficient for accurate conformational sampling, particularly in cases involving bond reorganization, while density functional theory (DFT), the standard for accurate energetics, is computationally prohibitive. Recent advances in semiempirical quantum methods (Froitzheim et al., 2025) and universal ML interatomic potentials (uMLIPs) demonstrate that near-DFT accuracy can be achieved at a fraction of the compute, and open the door to scalable and accurate sampling. To illustrate this direction and compare the computational cost in practical settings, we apply WT-ASBS to sample reactive energy landscapes involving bond formation (Trizio & Parrinello, 2021). We employ a recent uMLIP, UMA-S-1.1 (Wood et al., 2025), which accurately reproduces DFT energies, as the underlying energy model. For reaction sampling, we cross-check with WTMetaD results since sampling the transition region with unbiased samplers like MD is practically impossible (e.g., $\sim 10^{15}$ samples required for a 20 kcal/mol barrier at 300 K).
Figure 4: Sampling reactive landscapes. (a) SN2 reaction scheme and CVs. (b) PMF $F(d_{1},d_{2})$ for the SN2 reaction. (c) Cycloaddition reaction scheme with post-TS bifurcation. (d) Three bond-formation distances. (e) Contact CVs $c_{i}$ derived from distances $d_{i}$. (f,g) 1D $F(s_{1})$ and 2D $F(s_{1},s_{2})$ PMFs for the post-TS bifurcation.
@@
We study a prototypical SN2 reaction (Figure˜4a), where a chloride ion attacks the tetrahedral carbon and another chloride ion departs, leading to interconversion between states 1 $\leftrightarrow$ 2. We set two C–Cl bond distances as CVs and apply restraints to prevent chloride ions from diffusing away, ensuring the CV space remains bounded. Given the small system size, a single configuration (state 1) suffices for pretraining the sampler. Using WT-ASBS, we obtain the 2-D PMF along the two bond distances (Figure˜4b). The symmetric PMF agrees well with the WTMetaD result (Figure˜9), and both the location and free energy barrier of the transition state (TS) coincide with the results from standard saddle-point optimization.
@@
For simple reactions like SN2 with a single reaction channel between reactants and products, the process can be well characterized by identifying the TS and evaluating single-point energies, with harmonic approximations describing the local energy landscape around the reaction progress (Truhlar et al., 1996; Peters, 2017). However, many realistic challenges in chemistry involve branched or multimodal energy landscapes that cannot be captured by such approximations and instead requires explicit sampling. The reaction in Figure˜4c exemplifies this difficulty: reactant 3 passes through a single TS region that bifurcates into two products 4 and 5. Because the TS is not uniquely associated with a single product (i.e., the TS is ambimodal), characterizing this reaction requires sampling the full reactive landscape.
To construct CVs, we represent the three relevant C–C distances (Figure˜4d) as normalized contacts in Equation˜55, capturing bond formation progress (Figure˜4e). From these we define two CVs: $s_{1}=c_{1}+c_{2}+c_{3}$, which tracks overall bond formation (a soft bond count), and $s_{2}=c_{2}-c_{3}$, which distinguishes between products 4 and 5. Details on restraints and settings are provided in the Appendix˜D. Figure˜4f and g show the PMFs $F(s_{1})$ and $F(s_{1},s_{2})$, and similarly to SN2 example, the PMFs agree well with the WTMetaD result (2-D comparison in Figure˜10), providing accurate energetics for complex, multimodal reactive landscapes that dominate practical challenges in chemistry. Further discussion on reaction sampling tasks is provided in Appendix˜E.
Figure 5: Convergence of PMFs. PMF MAE from the final PMF for each reaction.
@@ 5 Discussion
We presented WT-ASBS, which integrates bias deposition over chosen CVs into a diffusion-based sampler, enabling exploration of unseen conformational modes and recovery of PMFs and importance sampling weights that reconstruct the Boltzmann distribution. The method unifies diffusion-based sampling with enhanced sampling and provides a practical and efficient recipe for sampling molecular systems, which allows diverse methodological extensions, including transition path sampling (Raja et al., 2025) and conformational exploration with system-agnostic CVs (e.g., RMSD).
@@
WT-ASBS requires the use of CVs, and while we utilized predefined ones, more complex systems may benefit from ML CVs and iterative cycles that combine exploratory WT-ASBS with CV training. A fair comparison of enhanced sampling methods is difficult because of the large number of hyperparameters involved (Hénin et al., 2022), and although WTMetaD may not be the optimal choice, we use it as a baseline since our enhanced sampling component is based on WTMetaD. Future work will explore combinations of different enhanced sampling strategies to provide a more comprehensive analysis. Finally, our approach does not allow amortization in the current form, while it is possible to construct the target bias variationally using neural networks (Bonati et al., 2019), which could be applied to systems that share similar slow degrees of freedom.
@@ A.1 Enhanced Sampling
One foundational approach is umbrella sampling. Introduced by Torrie & Valleau (1977), umbrella sampling runs simulations with an added bias potential $U(s)$ along a chosen reaction coordinate or collective variable $s=\xi(x)$. By choosing a series of overlapping bias “umbrellas” that cover the range of $\xi$, one can force the system to sample even high-energy configurations in each window. Each window’s samples are then reweighted by $\exp(+\beta U(s))$ to reconstruct the unbiased free energy landscape. This method demonstrated the principle of biasing along important coordinates to access otherwise inaccessible states. However, it requires a careful tuning of windows.
@@
Metadynamics took adaptive biasing a step further by dynamically building the bias potential during a single simulation. In the original metadynamics (Laio & Parrinello, 2002), the simulation deposits small Gaussian hills on the free energy landscape along the chosen CVs every few timesteps. These hills accumulate in wells (low free energy regions), gradually filling them and effectively raising the energy floor so the system escapes. Over time, the bias $V_{t}(s)$ approximates the negative of the underlying free energy F(s)-F(s), flattening the CV landscape. Well-tempered metadynamics (Barducci et al., 2008) refined this approach by gradually reducing the height of added hills as the bias grows. This means early in the simulation, the bias grows quickly (promoting exploration), but later it converges towards a finite bias that yields a modified stationary distribution rather than diverging. Specifically, well-tempering introduces a bias factor $\gamma>1$ so that the deposited bias is scaled by a factor that depends on the current bias (preventing overfill). The resulting steady-state is the well-tempered distribution in CV space, in which the free energy barriers are lowered by a factor of $\gamma$, effectively simulating the system at a higher temperature $T_{\text{eff}}=\gamma T$ along those CVs. Importantly, even with bias, samples can be reweighted to recover the true Boltzmann distribution. Well-tempered metadynamics thus balances exploration and accuracy by never completely flattening the landscape, avoiding infinitely large weights and ensuring the bias converges asymptotically to a finite correction.
@@
In the enhanced sampling community, there have been continual improvements to biasing algorithms. Notably, on-the-fly probability enhanced sampling (OPES) by Invernizzi & Parrinello (2020) is a performant metadynamics variant where the bias is constructed by directly targeting a pre-defined distribution (often the well-tempered distribution) at each step rather than depositing fixed-shape hills. This results in smoother bias updates and often faster convergence. Although we do not explicitly implement OPES, our algorithm likewise targets the well-tempered distribution Equation˜7. In a sense, WT-ASBS could incorporate OPES-style updates (adjusting the bias based on deviation between current CV histogram and target distribution), which could be an interesting direction for future work. Additionally, techniques like variationally optimized enhanced sampling (VES) (Valsson & Parrinello, 2014) cast bias finding as an optimization problem of $V(s)$ to minimize the KL divergence between biased and unbiased ensembles. Such approaches could also be combined with our generative sampler by using the model’s output to estimate that KL objective. In our current work, we use the simpler hill-based scheme inspired by WTMetaD.
@@
Beyond metadynamics, numerous enhanced sampling strategies exist. The adaptive biasing force (ABF) method (Darve et al., 2008) takes a different approach: instead of depositing hills, it continuously estimates the mean force along the CV and subtracts it, effectively flattening the free energy gradient. ABF and its extensions (e.g., extended ABF in Comer et al. (2015)) guarantee convergence by driving the observed force toward zero in well-sampled regions. In practice, ABF and metadynamics often achieve similar outcomes: a nearly uniform sampling of $s$, but via force-balancing vs. potential-filling mechanisms. Another family of methods is replica exchange and parallel tempering. In replica exchange MD (Sugita & Okamoto, 1999), multiple copies of the system are simulated at different temperatures (or Hamiltonians) in parallel. These replicas occasionally swap configurations, allowing a cold replica to receive a configuration from a hot replica that has overcome barriers. Over many swaps, each replica samples from its designated ensemble (e.g., the lowest-temperature one samples the target distribution) but benefits from exchanges to traverse phase space faster. Parallel tempering effectively raises the chance of barrier crossing by leveraging high-temperature dynamics, albeit at the computational cost of running many simulations. It is a powerful method to explore complex landscapes, though it requires careful selection of the temperature ladder and has high computational overhead if many replicas are needed.
@@ A.2 Neural Samplers
As opposed to MCMC and MD methods, generative models can potentially be used to generate samples from multimodal distributions without getting trapped in local minima or requiring lengthy equilibration times. Using generative models to sample probability distributions specified by a potential function was originally proposed by Noé et al. (2019); Albergo et al. (2019), and Nicoli et al. (2020) in the contexts of many-body physics, lattice field theory, and statistical mechanics, respectively. These initial approaches trained discrete-time normalizing flows (Tabak & Turner, 2013) by optimizing the KL divergence between the model and the unnormalized target densities. As the use of reverse-KL objective is prone to mode collapse, the community explored alternative training objectives for flow-based sampling (Vaitl et al., 2022; Midgley et al., 2022; Máté & Fleuret, 2023). However, using flow-based methods with these approaches presents challenges. Computing the likelihood of continuous flow models (Chen et al., 2018) is computationally expensive, yet it is a necessary component of many of these potential function-based learning objectives. Meanwhile, designing expressive discrete-time normalizing flows is complex due to the restrictions on their coupling layers (Dinh et al., 2016; Midgley et al., 2023).
Thinking of the sampling task as a problem of measure transport opens up other lines of research which make use of ideas introduced by denoising diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) and related, more general frameworks for generative modeling (Lipman et al., 2022; Albergo et al., 2023). In data-based settings, these approaches use intermediate distributions to transport samples to the target distribution, guided by a local quadratic regression (“matching”) objective that ensures the correct change of measure. The simulation-free nature of these approaches, combined with their straightforward training objectives, makes them more scalable and easier to train than previous methods, establishing them as the leading generative paradigm today. In the data-free setting, however, since samples are not available, trajectories must be sampled from the model. The central goal of recent works has thus been to design training objectives that (i) lead to a correct transport of measures and (ii) do not require backpropagation through the learned dynamics.
Vargas et al. (2023a); Berner et al. (2022) optimize the KL divergence in path space and establish connections between generative modeling and SOC. Zhang & Chen (2021); Havens et al. (2025) also approach the problem from the viewpoint of SOC by minimizing the control objective with the terminal cost defined by the target energy. Richter & Berner (2023) and Vargas et al. (2023b) optimize the variance of Radon–Nikodym derivative between the forward and backward path measures. Albergo & Vanden-Eijnden (2024) generalizes annealed importance sampling (AIS) by learning an SDE that minimizes the impact of unbiasing weights of AIS.
@@
Recent methods attempt to alleviate the challenge of energy-based training struggling to capture multiple modes in complex energy landscapes. Schopmans & Friederich (2025) applied temperature annealing in Boltzmann Generators, and concurrently Akhound-Sadegh et al. (2025); Rissanen et al. (2025) incorporated inference-time annealing into diffusion models to sample from Boltzmann densities. These approaches smoothly interpolate from a high-temperature (smoothed-out) version of the target distribution to the true low-temperature distribution akin to parallel tempering. Our method biases the sampler on the fly along specific CVs, which can be seen as a more targeted form of annealing that raises the effective sampling temperature only along selected important degrees of freedom.
@@
Another line of related work focuses on directly encouraging exploration during training. Kim et al. (2025) improved training efficiency by incorporating MCMC with bias from random network distillation during the search stage, steering exploration toward configurations that would otherwise be rarely visited, conceptually similar to our goal of amplifying exploration. In recent work, Blessing et al. (2025) formulate a diffusion-based sampler with a trust-region constrained optimal control perspective, which effectively adds Lagrange multiplier terms to enforce constraints on the path distribution. Our approach is unique in that it incorporates an adaptive bias during training that specifically targets low-dimensional CVs, which is highly effective for exploring rare-event regions and further enables reconstruction of the PMF.
A.3 Molecular Isomerism

### v2

@@ 4.2 Sampling Reactive Energy Landscapes
While peptide results highlight the efficiency of WT-ASBS in terms of the number of energy evaluations, the computational budgets in diffusion-based sampling in these cases are dominated by SDE integration with a control network, since classical force fields are extremely fast to evaluate. However, classical force fields are insufficient for accurate conformational sampling, particularly in cases involving bond reorganization, while density functional theory (DFT), the standard for accurate energetics, is computationally prohibitive. Recent advances in semiempirical quantum methods (froitzheim2025gxtb) and universal ML interatomic potentials (uMLIPs) demonstrate that near-DFT accuracy can be achieved at a fraction of the compute, and open the door to scalable and accurate sampling. To illustrate this direction and compare the computational cost in practical settings, we apply WT-ASBS to sample reactive energy landscapes involving bond formation (trizio2021enhanced). We employ a recent uMLIP, UMA-S-1.1 (wood2025uma), which accurately reproduces DFT energies, as the underlying energy model. For reaction sampling, we cross-check with WTMetaD results since sampling the transition region with unbiased samplers like MD is practically impossible (e.g., $\sim 10^{15}$ samples required for a 20 kcal/mol barrier at 300 K).
Figure 4: Sampling reactive landscapes. (a) SN2 reaction scheme with nucleophilic substitution between states 1 and 2, along with the two bond distance CVs $d_{1}$ and $d_{2}$ that describe approach and departure of the chloride ions. (b) PMF $F(d_{1},d_{2})$ for the SN2 system, showing the reactant and product basins and the transition-state region along the two-dimensional distance coordinates. (c) Cycloaddition reaction scheme exhibiting a post-TS bifurcation, highlighting the competing pathways that originate from the same transition structure. (d) Definition of the three forming bond distances $d_{1}$, $d_{2}$, and $d_{3}$ used to parameterize progress along the bifurcating reaction pathways. (e) Contact CVs $c_{i}$ obtained from the distances $d_{i}$, used to construct smoother CVs for sampling. (f) One-dimensional PMF $F(s_{1})$ along the CV $s_{1}=c_{1}+c_{2}+c_{3}$, comparing WT-ASBS and WTMetaD. (g) Two-dimensional PMF $F(s_{1},s_{2})$ with $s_{2}=c_{2}-c_{3}$, resolving the post-TS bifurcation and showing the distinct product channels, with TS locations from saddle point optimization.
@@
We study a prototypical SN2 reaction (Figure˜4a), where a chloride ion attacks the tetrahedral carbon and another chloride ion departs, leading to interconversion between states 1 $\leftrightarrow$ 2. We set two C–Cl bond distances as CVs and apply restraints to prevent chloride ions from diffusing away, ensuring the CV space remains bounded. Given the small system size, a single configuration (state 1) suffices for pretraining the sampler. Using WT-ASBS, we obtain the 2-D PMF along the two bond distances (Figure˜4b). The symmetric PMF agrees well with the WTMetaD result (Figure˜16), and both the location and free energy barrier of the transition state (TS) coincide with the results from standard saddle-point optimization.
@@
For simple reactions like SN2 with a single reaction channel between reactants and products, the process can be well characterized by identifying the TS and evaluating single-point energies, with harmonic approximations describing the local energy landscape around the reaction progress (truhlar1996current; peters2017reaction). However, many realistic challenges in chemistry involve branched or multimodal energy landscapes that cannot be captured by such approximations and instead requires explicit sampling. The reaction in Figure˜4c exemplifies this difficulty: reactant 3 passes through a single TS region that bifurcates into two products 4 and 5. Because the TS is not uniquely associated with a single product (i.e., the TS is ambimodal), characterizing this reaction requires sampling the full reactive landscape.
Figure 5: Convergence of PMFs. PMF MAE from the final converged PMF for each reaction and method, compared over wall-clock time.
To construct CVs, we represent the three relevant C–C distances (Figure˜4d) as normalized contacts in Equation˜55, capturing bond formation progress (Figure˜4e). From these we define two CVs: $s_{1}=c_{1}+c_{2}+c_{3}$, which tracks overall bond formation (a soft bond count), and $s_{2}=c_{2}-c_{3}$, which distinguishes between products 4 and 5. Details on restraints and settings are provided in the Appendix˜D. Figure˜4f and g show the PMFs $F(s_{1})$ and $F(s_{1},s_{2})$, and similarly to SN2 example, the PMFs agree well with the WTMetaD result (2-D comparison in Figure˜17), providing accurate energetics for complex, multimodal reactive landscapes that dominate practical challenges in chemistry. Further discussion on reaction sampling tasks is provided in Appendix˜F.
@@ 5 Discussion
We presented WT-ASBS, which integrates well-tempered bias deposition over chosen CVs into a diffusion-based sampler, enabling exploration of unseen conformational modes and recovering PMFs and importance sampling weights that reconstruct the Boltzmann distribution. By embedding enhanced sampling directly into the sampler, WT-ASBS links diffusion-based generative samplers with classical rare event techniques and provides a practical recipe for sampling molecular systems with controllable bias strength and straightforward statistical reweighting. We have also demonstrated that WT-ASBS can be combined with ML CVs (Section˜F.4), illustrating that the method is compatible with data-driven representations and can be used as a downstream sampler once informative low-dimensional coordinates have been learned. Looking ahead, we expect the most effective workflows to combine energy-based sampling with data-based generative models and to mix global moves (diffusion-based proposals in configuration or CV space) with local sampling (short MD or MCMC) in a problem-dependent way, for example, using global diffusion moves for slow solute rearrangements and local samplers for fast solvent motions. Finally, although this work focuses on sampling molecular configurations, low-dimensional latent variables have proven useful in discrete lattice sampling (swendsen1987nonuniversal; wang2001efficient) and dynamic Bayesian networks (doucet2000rao), suggesting that extending the WT-ASBS-like approach beyond molecular systems would be a fruitful direction.
@@
WT-ASBS requires the specification of CVs, and we primarily utilized predefined CVs and extended to ML CVs where metastable states are known. More complex systems will benefit from ML CVs and iterative cycles that interleave exploratory WT-ASBS runs with CV training updates. A fair comparison of enhanced sampling methods remains challenging due to the large number of algorithmic and system-specific hyperparameters (henin2022enhanced), and although WTMetaD may not be the optimal baseline for every system, we use it here as a natural baseline because our bias deposition scheme is inherited from WTMetaD. Future work will explore combinations of different enhanced sampling strategies to provide a more comprehensive analysis. Finally, WT-ASBS does not provide amortized sampling across systems or thermodynamic conditions, and extending it with a variationally parameterized bias (bonati2019neural) or shared control networks across related systems is a promising direction for transferring biases between systems that share similar slow modes.
@@ A.1 Enhanced Sampling
One foundational approach is umbrella sampling. Introduced by torrie1977nonphysical, umbrella sampling runs simulations with an added bias potential $U(s)$ along a chosen reaction coordinate or collective variable $s=\xi(x)$. By choosing a series of overlapping bias “umbrellas” that cover the range of $\xi$, one can force the system to sample even high-energy configurations in each window. Each window’s samples are then reweighted by $\exp(+\beta U(s))$ to reconstruct the unbiased free energy landscape. This method demonstrated the principle of biasing along important coordinates to access otherwise inaccessible states. However, it requires a careful tuning of windows.
@@
Metadynamics took adaptive biasing a step further by dynamically building the bias potential during a single simulation. In the original metadynamics (laio2002escaping), the simulation deposits small Gaussian hills on the free energy landscape along the chosen CVs every few timesteps. These hills accumulate in wells (low free energy regions), gradually filling them and effectively raising the energy floor so the system escapes. Over time, the bias $V_{t}(s)$ approximates the negative of the underlying free energy F(s)-F(s), flattening the CV landscape. Well-tempered metadynamics (barducci2008well) refined this approach by gradually reducing the height of added hills as the bias grows. This means early in the simulation, the bias grows quickly (promoting exploration), but later it converges towards a finite bias that yields a modified stationary distribution rather than diverging. Specifically, well-tempering introduces a bias factor $\gamma>1$ so that the deposited bias is scaled by a factor that depends on the current bias (preventing overfill). The resulting steady-state is the well-tempered distribution in CV space, in which the free energy barriers are lowered by a factor of $\gamma$, effectively simulating the system at a higher temperature $T_{\text{eff}}=\gamma T$ along those CVs. Importantly, even with bias, samples can be reweighted to recover the true Boltzmann distribution. Well-tempered metadynamics thus balances exploration and accuracy by never completely flattening the landscape, avoiding infinitely large weights and ensuring the bias converges asymptotically to a finite correction.
@@
In the enhanced sampling community, there have been continual improvements to biasing algorithms. Notably, on-the-fly probability enhanced sampling (OPES) by invernizzi2020rethinking is a performant metadynamics variant where the bias is constructed by directly targeting a pre-defined distribution (often the well-tempered distribution) at each step rather than depositing fixed-shape hills. This results in smoother bias updates and often faster convergence. Although we do not explicitly implement OPES, our algorithm likewise targets the well-tempered distribution Equation˜7. In a sense, WT-ASBS could incorporate OPES-style updates (adjusting the bias based on deviation between current CV histogram and target distribution), which could be an interesting direction for future work. Additionally, techniques like variationally optimized enhanced sampling (VES) (valsson2014variational) cast bias finding as an optimization problem of $V(s)$ to minimize the KL divergence between biased and unbiased ensembles. Such approaches could also be combined with our generative sampler by using the model’s output to estimate that KL objective. In our current work, we use the simpler hill-based scheme inspired by WTMetaD.
@@
Beyond metadynamics, numerous enhanced sampling strategies exist. The adaptive biasing force (ABF) method (darve2008adaptive) takes a different approach: instead of depositing hills, it continuously estimates the mean force along the CV and subtracts it, effectively flattening the free energy gradient. ABF and its extensions (e.g., extended ABF in comer2015adaptive) guarantee convergence by driving the observed force toward zero in well-sampled regions. In practice, ABF and metadynamics often achieve similar outcomes: a nearly uniform sampling of $s$, but via force-balancing vs. potential-filling mechanisms. Another family of methods is replica exchange and parallel tempering. In replica exchange MD (sugita1999replica), multiple copies of the system are simulated at different temperatures (or Hamiltonians) in parallel. These replicas occasionally swap configurations, allowing a cold replica to receive a configuration from a hot replica that has overcome barriers. Over many swaps, each replica samples from its designated ensemble (e.g., the lowest-temperature one samples the target distribution) but benefits from exchanges to traverse phase space faster. Parallel tempering effectively raises the chance of barrier crossing by leveraging high-temperature dynamics, albeit at the computational cost of running many simulations. It is a powerful method to explore complex landscapes, though it requires careful selection of the temperature ladder and has high computational overhead if many replicas are needed.
@@ A.2 Neural Samplers
As opposed to MCMC and MD methods, generative models can potentially be used to generate samples from multimodal distributions without getting trapped in local minima or requiring lengthy equilibration times. Using generative models to sample probability distributions specified by a potential function was originally proposed by noe2019boltzmann; albergo2019flow, and nicoli2020asymptotically in the contexts of many-body physics, lattice field theory, and statistical mechanics, respectively. These initial approaches trained discrete-time normalizing flows (tabak2013family) by optimizing the KL divergence between the model and the unnormalized target densities. As the use of reverse-KL objective is prone to mode collapse, the community explored alternative training objectives for flow-based sampling (vaitl2022gradients; midgley2022flow; mate2023learning). However, using flow-based methods with these approaches presents challenges. Computing the likelihood of continuous flow models (chen2018neural) is computationally expensive, yet it is a necessary component of many of these potential function-based learning objectives. Meanwhile, designing expressive discrete-time normalizing flows is complex due to the restrictions on their coupling layers (dinh2016density; midgley2023se).
Thinking of the sampling task as a problem of measure transport opens up other lines of research which make use of ideas introduced by denoising diffusion models (sohl2015deep; ho2020denoising) and related, more general frameworks for generative modeling (lipman2022flow; albergo2023stochastic). In data-based settings, these approaches use intermediate distributions to transport samples to the target distribution, guided by a local quadratic regression (“matching”) objective that ensures the correct change of measure. The simulation-free nature of these approaches, combined with their straightforward training objectives, makes them more scalable and easier to train than previous methods, establishing them as the leading generative paradigm today. In the data-free setting, however, since samples are not available, trajectories must be sampled from the model. The central goal of recent works has thus been to design training objectives that (i) lead to a correct transport of measures and (ii) do not require backpropagation through the learned dynamics.
vargas2023denoising; berner2022optimal optimize the KL divergence in path space and establish connections between generative modeling and SOC. zhang2021path; havens2025adjoint also approach the problem from the viewpoint of SOC by minimizing the control objective with the terminal cost defined by the target energy. richter2023improved and vargas2023transport optimize the variance of Radon–Nikodym derivative between the forward and backward path measures. albergo2024nets generalizes annealed importance sampling (AIS) by learning an SDE that minimizes the impact of unbiasing weights of AIS.
@@
Recent methods attempt to alleviate the challenge of energy-based training struggling to capture multiple modes in complex energy landscapes. schopmans2025temperatureannealed applied temperature annealing in Boltzmann Generators, and concurrently akhoundsadegh2025progressive; rissanen2025progressive incorporated inference-time annealing into diffusion models to sample from Boltzmann densities. These approaches smoothly interpolate from a high-temperature (smoothed-out) version of the target distribution to the true low-temperature distribution akin to parallel tempering. Our method biases the sampler on the fly along specific CVs, which can be seen as a more targeted form of annealing that raises the effective sampling temperature only along selected important degrees of freedom.
@@
Another line of related work focuses on directly encouraging exploration during training. kim2025scalable improved training efficiency by incorporating MCMC with bias from random network distillation during the search stage, steering exploration toward configurations that would otherwise be rarely visited, conceptually similar to our goal of amplifying exploration. In recent work, blessing2025trust formulate a diffusion-based sampler with a trust-region constrained optimal control perspective, which effectively adds Lagrange multiplier terms to enforce constraints on the path distribution. Our approach is unique in that it incorporates an adaptive bias during training that specifically targets low-dimensional CVs, which is highly effective for exploring rare-event regions and further enables reconstruction of the PMF.
A.3 Extended Related Works
ML-based transition path sampling
While our method is designed for equilibrium sampling with a CV-based biasing inside an energy-based diffusion sampler, closely related ideas have been developed for sampling rare dynamical trajectories by learning control forces on trajectory space, in the context of transition path sampling (TPS). holdijk2023stochastic cast molecular TPS as an SOC problem and learn CV-free policies that steer trajectories between metastable states. du2024doob propose Doob’s Lagrangian, a variational formulation of Doob’s $h$-transform (doob1957conditional) that learns conditional dynamics for generating transition paths between prescribed endpoints. seong2025transition introduce a diffusion-based TPS with off-policy training based on log-variance divergence between biased MD and target path measures. Different variational approaches to path sampling of rare events are reviewed in singh2025variational, with a focus on trajectory ensemble reweighting and optimal control approaches. Finally, raja2025action repurpose pretrained generative models for TPS by interpreting paths as trajectories of an induced SDE and finding most-probable transition paths via Onsager–Machlup action minimization.
Data-based diffusion sampling in chemistry
Complementary to our energy-based sampler that only assumes access to the potential energy and its gradients, several works train diffusion or flow-based generative models directly on molecular simulation or thermodynamic data to build data-driven equilibrium samplers. wang2022data train diffusion models on ensembles at multiple thermodynamic conditions to generate molecular configurations across temperature or parameter space. herron2024inferring develop thermodynamic maps, a diffusion-like generative approach to model the joint distribution of configurations and auxiliary thermodynamic variables to reconstruct phase behavior and other macroscopic properties from sparse data. qiu2025latent develop latent thermodynamic models that jointly learn low-dimensional representations and temperature-dependent equilibrium distributions. benayad2025hamiltonian combine replica exchange simulations with diffusion models and importance sampling to refine conformational ensembles and explore free-energy landscapes along chosen CVs.
A.4 Molecular Isomerism